OpenStudy (anonymous):

PLEASE HELP! An aircraft flew a distance of 3000 km from Berlin to Cairo at an average speed of v km/h. i) Write down an expression for time in hours what it took for the journey. ii)The aircraft returned non-stop from by the same route at an average speed of 2v km/h.write an expression to show this. iii)Write down an expression for the time in hours iv)Given that the time for the journey is 4 hours, form an equation in v and solve it. Guys i need you!

OpenStudy (anonymous):

OK so do you know the basic equation, rate * time = distance?

OpenStudy (anonymous):

Yes I do

OpenStudy (anonymous):

OK, so for part (i), plus in the rate and the distance, then isolate t.

OpenStudy (anonymous):

Is it: t=3000/v?

OpenStudy (anonymous):

yep, exactly

OpenStudy (anonymous):

Now the same for (ii)

OpenStudy (anonymous):

so 3000/2v?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Now you have written two expressions for the time out and the time back. So if you know the total time for the round trip is 4 hours, can you express that as an equation?

OpenStudy (anonymous):

uhm....4=3000/3v ?

OpenStudy (anonymous):

no, you can't combine unlike denominators. It's just 4 = (the sum of your two previous expressions)

OpenStudy (anonymous):

4=6000/v? This is where i start gettin a bit confused

OpenStudy (anonymous):

it's really simple, just literally put a plus sign between the two expressions you computed in part (i) and part (ii)

OpenStudy (anonymous):

Are you saying: 4=3000/v+3000/2v?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

just to be clear, it's \[4 = {3000 \over v} + {2000 \over 2v} \]

OpenStudy (anonymous):

Now to solve you multiply the whole thing through by 2v, which means multiplying each term by 2v.

OpenStudy (anonymous):

So 4*2v=2v*3000/v+2v*3000/2v

OpenStudy (anonymous):

right, now cancel out the extra v's, what do you get?

OpenStudy (anonymous):

you should get an equation with no division in it any more

OpenStudy (anonymous):

2v*4=v3000+3000

OpenStudy (anonymous):

exactly! now collect the v's on one side

OpenStudy (anonymous):

2v-3000v=3000/4

OpenStudy (anonymous):

hmm, something is fishy here, let me try and figure out what went wrong

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

ack, ok, haha

OpenStudy (anonymous):

So if you have 4 = (3000/v) + (3000/2v) and you multiply by v, all the v's on the right cancel. you get 4 * 2v = 2v*(3000/v) + 2v*(3000/2v)

OpenStudy (anonymous):

All the v's cancel out of the right side.

OpenStudy (anonymous):

I keep getting 2v*4=3000v+3000 as when the 2v is * by the 3000/v I would end up with that.

OpenStudy (anonymous):

By the way thanks for the help so far